jimmy is a partner in an internet based coffee supplier. The company offers gourmet coffee beans for $11 per pound and regular coffee beans for $7 per pound. Jummy is creating a medium price product that will sell for $9 per pound. The first thing to go into the mixing bin was 10 pounds of the gourmet beans. how many pounds of the less expensive regular beans should be added?
Answer by nerdybill(7384)

jimmy is a partner in an internet based coffee supplier. The company offers gourmet coffee beans for $11 per pound and regular coffee beans for $7 per pound. Jummy is creating a medium price product that will sell for $9 per pound. The first thing to go into the mixing bin was 10 pounds of the gourmet beans. how many pounds of the less expensive regular beans should be added?
.
Let x = amount (pounds) of lest expensive beans ($7 per pound)
then
10-x = amount (pounds) of $11 per pound beans
.
7x + 11(10-x) = 9(10)
7x + 110-11x = 90
110-4x = 90
-4x = -20
x = 5 pounds